Prove that (1 - sin 2x)/(cos 2x)=(cos 2x)/(1 + sin 2x)

Answer by Alan3354(21605)

Prove that (1 - sin 2x)/(cos 2x)=(cos 2x)/(1 + sin 2x)
------------
Cross multiply
(1-sin(2x))*(1+sin(2x)) = cos^2(2x)
1 - sin^2(2x) = cos^2(2x) That's an identity.